OpenStudy (anonymous):

Explain how you would find the product of 2012 and 208. Include either the exponent rule or the expanded method in your explanation. Give your answer in exponential form

Parth (parthkohli):

\[\rm\large 2012 \times208 = (2000+12)(200 +8) \]agree?

OpenStudy (anonymous):

opps its suppose to be 20^12 and 20^8

Parth (parthkohli):

Oh, then it's a very famous rule.\[\rm \large a^m \times a^n = a^{m +n}\]

Parth (parthkohli):

Can you manage the rest?

OpenStudy (anonymous):

Not really but ill try

OpenStudy (anonymous):

A kilobyte is 2^10. A megabyte is 2^20 bytes. How many kilobytes are in a megabyte? Use the exponent rule for division to find your answer and give your answer in exponential form.